there is a similar question on here but it doesnt show how to solve it.
The perimeter of a rectangle yard is 270 feet. If its length is 25 feet greater than its width what are the dimensions of the yard?

Answer by checkley77(12844)

L=W+25
2L+2W=270
2(W+25)+2W=270
2W+50+2W=270
4W=270-50
4W=220
W=220/4
W=55 THE WIDTH.
L=55+25=80 THE LENGTH.
PROOF:
2*80+2^55=270
160+110=270
270=270